# Werewolf: A Straightforward Game Framework with TTS for Improved User Engagement

### v1

@@ Abstract
The growing popularity of social deduction game systems for both business applications and AI research has greatly benefited from the rapid advancements in Large Language Models (LLMs), which now demonstrate stronger reasoning and persuasion capabilities. Especially with the raise of DeepSeek R1 and V3 models, LLMs should enable a more engaging experience for human players in LLM-agent-based social deduction games like Werewolf. Previous works either fine-tuning, advanced prompting engineering, or additional experience pool to achieve engaging text-format Werewolf game experience. We propose a novel yet straightforward LLM-based Werewolf game system with tuned Text-to-Speech(TTS) models designed for enhanced compatibility with various LLM models, and improved user engagement. We argue with ever enhancing LLM reasoning, extra components will be unnecessary in the case of Werewolf.
Keywords Social Deduction Game Human-AI Interaction Human-Computer Interaction
@@ 1 Introduction
The Werewolf game, a widely recognized social deduction game, continues to attract increasing attention from AI researchers, particularly within the natural language processing (NLP) community. Prior work has introduced dedicated datasets for this domain [1] and leveraged them to fine-tune models for experience retrieval, enhancing large language model (LLM) performance in gameplay [2, 3]. These approaches often integrate external modules to support decision-making, but we argue that sufficiently advanced LLMs—such as DeepSeek V3, DeepSeek R1, and ChatGPT 4o—should be capable of playing the Werewolf game independently, given robust reasoning capabilities. While techniques like Chain-of-Thought (CoT) prompting can further enhance an LLM’s deductive reasoning by encouraging step-by-step thinking, they also introduce additional latency. Therefore, minimizing reliance on both external modules and verbose prompting strategies is essential for deploying responsive, real-time game systems where interaction speed is critical.
With the continuously improving reasoning capabilities of large language models (LLMs), recent studies have demonstrated that fine-tuned or state-of-the-art LLMs can achieve competitive performance in poker game scenarios [4, 5]. We anticipate that, as LLMs become more proficient in reasoning and in-context learning, fine-tuning and advanced prompting techniques will become increasingly unnecessary across many strategic settings. Motivated by this trend, we propose a straightforward Werewolf game framework in which we provide each simulated player with relevant context, such as game rules and general strategies, directly through prompts. The LLMs are then tasked with role-playing distinct player personas, strategically interacting to create an engaging and dynamic game experience.
In addition, the verbal communication aspect of the Werewolf game as an essential feature of its social dynamics has been largely overlooked in existing frameworks [2, 3]. To address this, we fine-tuned and integrated a text-to-speech (TTS) module into our system, enabling gameplay in both textual and auditory formats. Importantly, while audio output enhances immersion, all core reasoning remains text-based.
2 Frameworks
Our proposed Werewolf game system is composed of the following modules:
Werewolf Game: The game pipeline that manages the game flow, executes player’s actions, and judge the game status.
Players: The human or LLM role-played game players, agents with varying action space depending on their assigned roles.
User Engagement: A pipeline that parallels LLM and TTS generations.
These modules operate in parallel to create a coherent, vocal and engaging Werewolf game experience, enabling smooth interaction between game agents and users while maintaining the natural dynamics of a social deduction game.
2.1 Werewolf Game
2.1.1 Players
We define human and AI players, where AI players are role-played by LLM. AI players’ names are prepared in order to demonstrate the TTS avatars we tuned, including Jay Chow, Jack Ma, Ma Baoguo, Yang Mi, Luo Xiang, Shen Yi, and Donald Trump. The prepared names are either famous Chinese businessman, actors, influencers, or famous politician. This design choice heavily relies on the outcome of TTS module. Meanwhile, human players have the automacy to enter their preferred names in the game. The game requires at least players to play.
2.1.2 Roles
Only roles are available for the demonstration purposes since the complexities continue exploding when introducing more roles. In this simplest Werewolf game setup, the action space for each roles are notated as:
Villager:
Werewolf:
Seer:
Witch:
Here, vote refers to the act of selecting a player’s name for elimination after the discussion phase of the game. Reason denotes the verbal exchange during which players debate and identify potentially suspicious individuals, typically those suspected of being werewolves, thereby supporting their voting decisions. Reveal represents the special ability of the Seer, who, during the night phase, can uncover the hidden role of one player per round. Intuitively, kill indicates the target selected by the assigned Werewolves for elimination during the night phase of each round.
We further define two critical resources associated with the Witch’s actions: , denoting the number of remaining cures available to save a player targeted by the Werewolves, and , indicating the number of poisons that can be used to eliminate a player during the night. The Witch’s night action requires two sequential decisions formatted as: Cure/None, NameToEliminate/None.
During the day phase, active players are required to reason collaboratively through discussion, followed by a voting phase in which they vote to eliminate the most suspicious player based on inconsistencies or contradictions observed during the conversation. In the night phase, active players with special roles may perform their designated actions in a predefined sequence.
2.2 Game Flow
We encapsulate a single gameplay of the Werewolf game into a class, where player roles are automatically assigned and essential components, such as the game log, are initialized by default. Within this class, the game’s continuation status is managed by the game.judge function, which determines whether the game should proceed or end based on whether the number of active werewolves significantly exceeds that of the remaining active players. The framework evaluates the game status twice per full cycle (i.e., one night phase and one day phase): once immediately after the night phase, and again after the day phase. During the night phase, actions are executed in a predefined order: werewolves (kill) → seer (reveal) → witch (night). In the day phase, the discussion stage comes first, where all active players reason about who the werewolves might be based on prior interactions. This is followed by a voting stage where each player votes to eliminate the player they find most suspicious. We summarize the streamlined game flow as follows.
2.3 LLM Role-play
Unlike traditional agentic system, LLM-powered agent systems excels at various aspects from dynamically following and executing the instructions, to interact with the environment that requires solid reasoning capabilities [6]. The capabilities of such LLM agents largely rely on the the capabilities of LLMs the system integrated [2]. We integrated one of the most state-of-the-art LLMs DeepSeek V3 compromising longer response time of DeepSeek R1 which is designed for reasoning-intensive tasks, and higher costs of OpenAI models. However, we still integrated the portals for other LLMs from HuggingFace, llama cpp, and OpenAI for future experiment and development.
2.3.1 Variables
We list the below variables for LLMs to play the Werewolf game for solid reasoning, and only when they are necessary to their assigned roles:
active_players: Players who are currently active (i.e., active status in the game).
werewolves: List of players who are assigned with Werewolf roles.
seer_dict: The seer’s privately revealed information about other players’ roles, only visible to the seer.
werewolf_log: The action history of the werewolves(e.g. what did they do previously in the night phase), visible only to the werewolves.
witch_log: The action history of the witch, visible only to the witch.
log: The public game history, including eliminations, votes, and discussions, visible to all players.
2.3.2 Prompt Design
We assume the LLMs are as capable to play the Werewolf game without additional complex design such as Chain-of-Thoughts(CoT) prompting or ReAct prompting, considering DeepSeek V3 is distilled from DeepSeek R1[7]. We simply prompt the LLM by providing the information of players to be role-played, description of the action space, and guidance of the reply format, and very little guidance about the game rules. Below is one example of the prompt design for discussion phase.
2.4 Text-to-Speech Module
The Text-to-Speech (TTS) module implemented in this framework leverages the GPT-SoVITS[8], a state-of-the-art model combining Generative Pre-trained Transformer (GPT) techniques with Speaker-adaptive voice conversion (SoVITS) technology. This module enables high-quality voice synthesis and cloning from minimal audio input only in one shot, significantly reducing the challenge of collecting training data.
Key features of the implemented TTS module include:
Rapid Training: Only approximately one minute of speech data is required to achieve effective fine-tuning, allowing rapid deployment.
Zero-shot Capability: With as little as five seconds of speaker audio, the module can perform robust text-to-speech conversion, enabling quick adaptation to new voice profiles.
Cross-Language Support: The module supports multilingual synthesis, accommodating Chinese, English, Japanese, and other languages, enhancing versatility for diverse linguistic applications. Currently we have only set up English and Chinese modes and fine-tuned the voice packs targeting Jay Chow, Jack Ma, Ma Baoguo, Yang Mi, Luo Xiang, Shen Yi, and Donald Trump.
The TTS module was fine-tuned specifically for our framework to (1) optimize voice quality and (2) bring familiar voices into the game, enhancing the overall player experience. The streamlined training and inference pipeline significantly improves both efficiency and flexibility, enabling targeted and customizable voice synthesis for a variety of use cases.
2.5 User Engagement
2.5.1 Latency
We take into account that the process of TTS generating speech takes a certain amount of time. Instead of using a completely linear process (LLM generates a piece of text and then enters the TTS function for audio output, and then proceeds to the next stage of LLM text generation after all the audio output is completed), we processed LLM and TTS generation in parallel. The LLM generation operates in flush mode, where tokens are produced incrementally rather than waiting for the entire output to be generated. The system continuously listens to the generated text and triggers the TTS model to produce audio output either at clause boundaries, identified by commas, or at full sentence boundaries, identified by periods. The resulting audio segments are managed by a queue that ensures they are played sequentially, maintaining the natural flow of speech.
The advantage of this is that when presenting the game in audio format, this processing pipeline significantly reduces perceived response latency by allowing speech to begin as soon as the first tokens are generated, rather than waiting for the full output. This enables more natural and immersive interactions, closely mimicking human conversation patterns which is crucial in a social deduction setting where timing and tone matter. By chunking speech at clause or sentence boundaries, the system enables more natural and fluid dialogue. The audio queue allows smoother turn-taking, potential interruptions, and scalability in multi-agent settings. This design also helps keep computational resources efficiently utilized.
2.5.2 Scalability
We set aside a parameter to select the number of human players, which brings higher scalability to the program, and allows us to select one or more human users to play the game with other AI players, which also fits the real-life scenario where the user wants to play the game by himself or with 1-2 friends, but the number of people is too limited to carry out the full game flow.
We have also kept the prompt and rules definitions for the various game characters separately, which also creates the possibility of adding more game characters (such as hunters, guards, wolf lords, etc.) at a later date.
3 Qualitative Results
@@ 4 Limitations
Second, the current implementation supports only a subset of standard Werewolf roles (e.g., Seer, Witch, Werewolf, Villager). Expanding the system to include more complex roles (e.g., Hunter, Cupid, Guard) would introduce deeper interdependencies and increase the difficulty of role alignment, offering a stronger test of agent adaptability.
@@
Fourth, while the system supports audio output via TTS for immersive interaction, it currently lacks voice input. This limits its usability in real-time spoken interactions and reduces the naturalness of gameplay in multiplayer or human-in-the-loop settings.

### v2

# Verbal Werewolf: Engage Users with Verbalized Agentic Werewolf Game Framework

@@ Abstract
The growing popularity of social deduction games has created an increasing need for intelligent frameworks where humans can collaborate with AI agents, particularly in post-pandemic contexts with heightened psychological and social pressures. Social deduction games like Werewolf, traditionally played through verbal communication, present an ideal application for Large Language Models (LLMs) given their advanced reasoning and conversational capabilities. Prior studies have shown that LLMs can outperform humans in Werewolf games, but their reliance on external modules introduces latency that left their contribution in academic domain only, and omit such game should be user-facing. We propose Verbal Werewolf, a novel LLM-based Werewolf game system that optimizes two parallel pipelines: gameplay powered by state-of-the-art LLMs and a fine-tuned Text-to-Speech (TTS) module that brings text output to life. Our system operates in near real-time without external decision-making modules, leveraging the enhanced reasoning capabilities of modern LLMs like DeepSeek V3 to create a more engaging and anthropomorphic gaming experience that significantly improves user engagement compared to existing text-only frameworks.
Keywords Social Deduction Game AI for Games Human-AI Interaction Human-Computer Interaction
@@ 1 Introduction
There is an increasing need for intelligent social deduction game frameworks which one or few people can play with collaboration with AI due to the post-pandemic psychological pressure and heavier social burden. Social deduction games, Werewolf as one of the most typical ones, are often played in forms of verbal communication, which we argue the large language models(LLMs) should be excelled at[1]. Previous works have already confirmed that LLMs have certain probability to defeat human players in the werewolf games [2] with additional model for estimating the probability of winning the game. Other further works similarly have also been shown that high-performance LLMs with external modules for strategy optimizations can play the werewolf games strategically without fine-tuning the language models [3, 4]. However, these related works are heavily dependent on external modules for decision-makings and often demonstrate the game in simple language only. In addition, external modules often bring excessive latency, and disable the anthropopathic type-writing flush response that only constrains their framework at an academic level, overlooked the low-delay and anthropopathic nature as an engaging social deduction game.
In light of this, we propose Verbal Werewolf to further explore a more engaging werewolf game framework by optimizing two pipelines, one for the gameplay powered by the most state-of-the-art LLMs, and the other one for our fine-tuned Text to Speech(TTS) module that brings plain text output to live. We demonstrate the Verbal Werewolf progresses both pipeline in parallel and in nearly real-time, improving the user experience drastically.
2 Implementation
We demonstrate the Verbal Werewolf framework in mainly two pipelines: Game Play that manages the game flow, such roles assignments, LLM queries, and TTS Module that transforms sentence-by-sentence text output to designated audio outputs and manage the aduio play flow.
2.1 Game Play
Game Flow We present the werewolf game in the simplest setting[5] involving only werewolf, villager, seer, and witch roles. The Game Play mainly manages the whole game flow by initially assign different roles randomly, and we mainly divide one round of play into daytime, when they should reason then vote who are more suspicious being werewolves, and nighttime, when different players executes their assigned actions per their roles in the sequence of {werewolves-seer-witch} We formally define their action space as follows:
| Role | Daytime Actions | Nighttime Actions | Props |
| Villager | reason, vote | ||
| Werewolf | kill | ||
| Seer | reveal | ||
| Witch | night | cures, poisons |
In more detail, werewolves each round can kill a specific active player when this action is called, with the system selecting a player from their available choices if they differ. The Seer is able to reveal the assigned role of a particular player each round. The Witch can additionally cure the player that werewolves attempt to eliminate each round if the witch still possesses cures, or poison a particular player of their choice if the witch still has poisons available — all within one concatenated night action.
AI Players LLM-powered agent systems excel at various aspects, from dynamically following and executing instructions to interacting with environments that require solid reasoning capabilities. Their performance relies on the capabilities of the integrated LLMs. We assume that nowadays LLMs are capable of playing the Werewolf or a similar form of game without additional complex design such as Chain-of-Thoughts (CoT) and ReAct prompting, or external experience pool[6].
We integrated DeepSeek V3, one of the most state-of-the-art LLMs, as it offers a practical balance that provides strong reasoning capabilities, one is distilled from the reasoning-focused DeepSeek R1, while avoiding the longer response times of DeepSeek R1 and the higher costs of OpenAI models. We have also integrated portals for other LLMs available from HuggingFace, llama.cpp, and OpenAI to facilitate future experiments and development.
Our approach is straightforward: we prompt the LLM by providing information about the players to be role-played, a description of the action space, guidance on reply format, and minimal guidance about the game rules.
2.2 Text-to-Speech(TTS) Module
We fine-tuned GPT-SoVITS [7], a state-of-the-art TTS model that already excels at fast one-shot voice cloning from brief audio samples (5-10 seconds) with universal weights, by creating specialized model weights for each of 8 celebrities. While the original model provides excellent general-purpose voice synthesis, our fine-tuned versions achieve superior fidelity in reproducing each celebrity’s unique vocal characteristics, bringing their familiar voices into the game with enhanced quality and improving the overall player experience.
Combining Game Flow and TTS modules, our agentic system employs parallel processing threads to achieve real-time voice interaction. As the Game Play generates AI player responses token by token, the TTS thread concurrently processes every half sentence, synthesizing audio and queuing playback. This streaming architecture eliminates traditional wait-for-completion bottlenecks brought by both LLMs and TTS models, allowing AI players to begin speaking while still formulating their complete responses, creating natural conversational flow that mirrors human gameplay dynamics.
3 Qualitative Analysis
@@ 4 Limitations
Second, the current implementation supports only a subset of standard Werewolf roles (e.g., Seer, Witch, Werewolf, Villager). Expanding the system to include more complex roles (e.g., Hunter, Cupid, Guard) would introduce deeper inter-dependencies and increase the difficulty of role alignment, offering a stronger test of agent adaptability.
@@ 5 Conclusion
Supplementary Materials
This section provides additional implementation details of the Verbal Werewolf, the variables provided to large language model (LLM) agents for role-play, and the prompt design used in different game phases.
Game Flow Implementation
Below algorithm describes the procedural logic of the Werewolf game, including initialization, role-specific actions during the night phase, and decision-making during the day phase. This step-by-step representation ensures that all players’ actions, role constraints, and phase transitions are systematically handled.
Variables for LLM Role-play
Table 1 summarizes the key in-game variables accessible to LLM agents. Access to each variable is restricted based on the player’s role that enables role-specific reasoning and alignment with real-world game constraints.
| Variable | Description | Visibility |
| active_players | List of all currently active players in the game. | All roles |
| werewolves | Identities of all players assigned the Werewolf role. | Werewolf only |
| seer_dict | Seer’s private revelations of other players’ roles. | Seer only |
| werewolf_log | History of Werewolf kill actions from previous nights. | Werewolf only |
| witch_log | History of the Witch’s actions during the night phase. | Witch only |
| log | Public game history including eliminations, votes, and discussions. | All roles |
Prompt Design for LLM Agents
Below illustrates an example prompt used in the discussion phase. This minimalistic design avoids complex prompting strategies such as CoT or ReAct, relying on the base reasoning capabilities of DeepSeek V3 [8]. The prompt provides role-specific information, available actions, and conversational history.